5. Calculate a 95% confidence interval for the mean assuming the values come from a normal population: 12, 16, 21, 27, 29, 33.
Answer #1

Solution, Given that Using calculator, x = 21.33 = 6,47044563 S is C = 95% = 0.95 At 95% confidence interval t x = 1-C = 1-0.